# High-Capacity and Low-PAPR BICM-OFDM Systems Using Non-Equiprobable and Non-Uniform Constellation Shaping With Clipping and Filtering

### v1

@@ IV-B OFDM Receiver With CNC
At the receiver, we consider performing CNC to mitigate the in-band distortion caused by CAF prior to channel decoding. The block diagram of the OFDM receiver with CNC is illustrated in Fig. 8. In the case of an AWGN channel, the received symbol corresponding to the $k$th subcarrier can be expressed as
$\tilde{Z}_{k}=\alpha Z_{k}+D_{k}+W_{k},$ (19)
where $D_{k}\in{\mathbb{C}}$ denotes the distortion component caused by the CAF process and $W_{k}\in{\mathbb{C}}$ represents the Gaussian noise with $W_{k}\sim\mathcal{CN}\left(0,N_{0}\right)$, both corresponding to the $k$th subcarrier. Assuming that the baseband OFDM signal follows a complex Gaussian distribution, the attenuation factor $\alpha$ due to clipping can be expressed as [13]
@@
$Y_{k}=H_{k}\left(\alpha Z_{k}+D_{k}\right)+W_{k},$ (21)
@@
$\tilde{Y}_{k}\triangleq\frac{Y_{k}}{H_{k}}=\alpha Z_{k}+D_{k}+W_{k}^{\prime},$ (22)
@@
    Apply the hard decision based on the constellation set $\mathcal{A}({M,N,\rho})$ to the equalized and normalized received signal $Y_{k}^{(K)}/\alpha$ (with $Y_{k}^{(0)}:=\tilde{Y}_{k}$) to estimate the transmitted symbol $\hat{Z}^{(K)}_{k}={\hat{X}_{I,k}^{(K)}}+j{\hat{X}_{Q,k}^{(K)}}$, which is given by
    $\begin{cases}{\hat{X}}_{I,k}^{(K)}=\arg\min_{{X}\in\mathcal{A}({M,N,\rho})}\left({X}-\Re\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2},\\ {\hat{X}}_{Q,k}^{(K)}=\arg\min_{{X}\in\mathcal{A}({M,N,\rho})}\left({X}-\Im\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2}.\end{cases}$ (23)
@@
    Apply the same CAF process at the transmitter to obtain the estimated in-band component, which can be modeled as
    $\hat{{U}}_{k}^{(K)}=\alpha\hat{{Z}}_{k}^{(K)}+\hat{{D}}_{k}^{(K)}.$ (24)
@@
    $=\alpha{Z}_{k}+\left({D}_{k}-\hat{{D}}_{k}^{(K)}\right)+{W}_{k}.$ (26)
@@
Note that since our NENU constellations are designed based on a one-dimensional constellation, the complexity associated with the hard decision in (23) is significantly lower than that of the two-dimensional constellation, similar to the argument of demapper complexity. This observation also justifies the use of one-dimensional constellations for our BICM-OFDM systems involving CAF and CNC [^5] It is possible to design CNC that takes into account the channel decoder output as well as re-encoder in its loop, but this approach makes not only the receiver complexity and latency prohibitive, but also the appropriate system design more challenging. .
@@ IV-C1 System Setup
In the following numerical results, we consider the specific system where the number of subcarriers is $N_{c}=1024$ and each subcarrier is modulated by $32$-PAM per dimension ($1024$-QAM per subcarrier symbol) with $R_{c}=5/6$. The resulting spectral efficiency is $4.166$ bits per dimension, or $8.333$ bits per complex channel use. Based on the preliminary analysis, the clipping ratio of $\gamma_{\text{CR}}=1.5$ is found to be the minimum acceptable value that does not impose malfunctioning of CNC operation666 Optimal clipping ratio may be mathematically analyzed in terms of signal-to-noise plus distortion power ratio (SNDR) when CAF is applied [45], but since our system makes use of CNC at the receiver, its error rate performance depends on several parameters such as the number of subcarriers and constellations [15]. Therefore, it may be mathematically intractable. . Although some residual clipping noise still remains as a result of CNC, the proposed system can be designed to operate properly with $\gamma_{\text{CR}}=1.5$. Both AWGN and fading channels are considered, where in the case of fading channels we adopt frequency-selective Rayleigh fading channel with an uncorrelated four-path equal-power delay profile as a specific example. The proposed NENU constellation used in each system is optimized based on the corresponding channel condition.
@@ IV-C2 Parameter Optimization
We optimize the system based on (10), where the BMI is computed using (8) by observing the received symbol that serves as an input to the channel decoder. Note that for the systems employing CAF and CNC, the channel becomes a mixture of AWGN and residual clipping noise, which may not be described by a tractable mathematical model. Therefore, we compute the resulting BMI based on simulations, making the optimization process more computationally intensive compared to the AWGN channels without CAF.
We consider the following three cases:
  1. A.
    Without CAF, both AWGN and fading channels
  2. B.
    With CAF and CNC, AWGN channel
  3. C.
    With CAF and CNC, fading channel
For each case, the following optimal parameter sets are found through preliminary simulations:
  1. A.
    $N=32$, $\rho=0.88$  (SNR = $26$ dB)
  2. B.
    $N=24$, $\rho=0.93$  (SNR = $27$ dB)
  3. C.
    $N=28$, $\rho=0.97$  (SNR = $33$ dB)
In what follows, we refer to the NENU constellations designed by the above parameter sets as Modes A, B, and C, respectively.
From the above results, we observe that the optimal number of constellation points depends on the channel conditions, which agrees with the trend discussed in Section III. Without CAF, $N=32$ is selected as optimal due to the high spectral efficiency assumption with $R_{c}=5/6$. In contrast, with CAF and CNC, smaller values of $N$ are preferred to better accommodate the residual nonlinear distortion that reduces the effective SNR.
@@
We note that constellation design based on exhaustive numerical optimization algorithms, such as particle swarm optimization (PSO), does not rely on a predefined set of constellation candidates. Consequently, optimizing the objective functions such as the BMI after CAF and CNC processing becomes computationally demanding. On the other hand, the proposed design requires evaluation over a limited number of constellation candidates determined only by the two parameters. It thus reduces the computational burden significantly and enables optimization in an efficient manner, even under such complicated system constraints.
IV-C3 Total BMI
@@
Figure 9: Comparison of BMI values for 32-PAM constellations over an AWGN channel. The following three systems are evaluated: A1) Without CAF, A2) with CAF only, and A3) with CAF/CNC. (Parameters: $\gamma_{\text{CR}}=1.5$ for CAF and $K_{\text{max}}=10$ for CNC.)
@@
Figure 10: Comparison of BMI values for 32-PAM constellations over a frequency-selective Rayleigh fading channel. The following two systems are evaluated: F1) Without CAF and F2) with CAF and CNC. (Parameters: $\gamma_{\text{CR}}=1.5$ for CAF and $K_{\text{max}}=10$ for CNC.)
@@
We evaluate the total BMI over an AWGN channel for the following three systems:
@@
    BICM-OFDM without CAF (no PAPR reduction)
@@
    BICM-OFDM with CAF but without CNC
@@
The results are compared in Fig. 9, where the parameters of the NENU constellations are fixed as either Mode A or B over the entire channel SNR. The uniform PAM is also shown as a reference. In A1, Mode A demonstrates superior performance compared to the others at the target information rate. On the other hand, in A2, the achievable rates of all constellations significantly deviate from the target as a result of severe clipping noise. Finally, in A3, the information rate is successfully recovered by CNC, allowing the target information rate to be achieved with a lower PAPR compared to A1. Mode B demonstrates superior performance, yielding a greater gain over Mode A. The results thus demonstrate the effectiveness of the proposed constellation design tailored to the BICM-OFDM system with CAF and CNC.
We next consider the frequency-selective Rayleigh fading channel for the following two systems:
@@
    BICM-OFDM without CAF (no PAPR reduction)
@@
The BMI values are compared in Fig. 10, where the parameters of the NENU constellations are fixed as either Mode A or C over the entire channel SNR, along with the uniform PAM as a reference. It is observed that for F1, the shaping gains achieved by the proposed constellations decrease compared to the results over the AWGN channel. This may be due to the fact that a single modulation scheme is applied across all subcarriers, while they experience different effective SNRs. Nevertheless, in F2, significant gains by the proposed shaping are still observed when compared at the target information rate. Additionally, Mode C (optimized for this system over fading channel) offers enhanced performance over Mode A.
@@ V Simulation Results
We evaluate the system-level performance of the BICM-OFDM systems employing the proposed NENU constellations in the presence of CAF and CNC. As performance measures, we consider the distribution of PAPR as well as BER. For the binary channel code, we utilize the original rate-$1/3$ turbo code with rate-compatible puncturing [46]. The simulation setup is summarized in TABLE II. We compare the three systems:
@@
Note that S3 is equivalent to a single-carrier system with rectangular pulse shaping filter and thus its PAPR has direct dependence on the constellation [47].
For each system, we evaluate the uniform PAM, the proposed NENU constellations (parameter sets: Modes A, B, and C), and the one-dimensional NUC of ATSC 3.0. For S3, the conventional minimum mean square error (MMSE) detector is implemented at the receiver for the frequency-selective fading channel.
TABLE II: Simulation setup.
OFDM 1024 subcarriers
Modulation 32-PAM per dimension
Coding Punctured turbo code
Decoding Log-MAP (iteration: 15)
Code rate 5/6
Codeword length 30720
Spectral efficiency $8.333$ bits per subcarrier
PAPR reduction CAF ($\gamma_{\text{CR}}=1.5$, $J=4$) or DFT precoding
PAPR evaluation Four-times oversampling
CNC $K_{\text{max}}=10$
Channel AWGN or Rayleigh fading
Fading channel 4-path equal-power delay profile
@@ V-A CCDF of PAPR
Figure 11: Comparison of the CCDF of PAPR among the three systems: conventional OFDM without CAF, OFDM with CAF, and OFDM with DFT precoding (single-carrier). For each system, the three PAM constellations (uniform, Mode A, and Mode B) are evaluated.
The CCDF curves of the PAPR among the three systems are compared in Fig. 11, where the three PAM constellations (uniform as well as NENU with Modes A and B) are compared.
We observe no difference among the three curves plotted for S1 and S2, suggesting that each subcarrier constellation does not affect PAPR distribution for OFDM. On the other hand, in S3, the selection of constellations significantly affects the dynamic range of the transmit signals, indicating that the use of constellation shaping increases the PAPR as expected [48], even though it is still less than the conventional OFDM (S1). In contrast, from the results of S2, we observe that CAF effectively reduces the peak power of OFDM signals, even in the presence of constellation shaping, achieving a PAPR reduction of $5.0$ dB compared to S1 when measured at a CCDF of $10^{-4}$.

### v2

@@ IV-B OFDM Receiver With CNC
At the receiver, we consider performing CNC to mitigate the in-band distortion caused by CAF prior to channel decoding [^5] It is possible to design CNC that takes into account the channel decoder output as well as re-encoder in its loop, but this approach makes not only the receiver complexity and latency prohibitive, but also the appropriate system design more challenging. . The block diagram of the OFDM receiver with CNC is illustrated in Fig. 8. In the case of an AWGN channel, the received symbol corresponding to the $k$th subcarrier can be expressed as
$\tilde{Y}_{k}=\alpha X_{k}+D_{k}+W_{k},$ (21)
where $D_{k}\in{\mathbb{C}}$ denotes the distortion component caused by the CAF process and $W_{k}\in{\mathbb{C}}$ represents the Gaussian noise with $W_{k}\sim\mathcal{CN}\left(0,N_{0}\right)$, both corresponding to the $k$th subcarrier. Assuming that the baseband OFDM signal follows a complex Gaussian distribution, the attenuation factor $\alpha$ due to clipping can be expressed as [ochiai2002performance]
@@
$Y_{k}=H_{k}\left(\alpha X_{k}+D_{k}\right)+W_{k},$ (23)
@@
$\tilde{Y}_{k}\triangleq\frac{Y_{k}}{H_{k}}=\alpha X_{k}+D_{k}+W_{k}^{\prime},$ (24)
@@
    Apply the hard decision based on the constellation set $\mathcal{A}({M,N,\rho})$ to the equalized and normalized received signal $Y_{k}^{(K)}/\alpha$ (with $Y_{k}^{(0)}:=\tilde{Y}_{k}$) to estimate the transmitted symbol $\hat{X}^{(K)}_{k}={\hat{A}_{I,k}^{(K)}}+j{\hat{A}_{Q,k}^{(K)}}$, which is given by
    $\begin{cases}{\hat{A}}_{I,k}^{(K)}=\arg\min_{{A}\in\mathcal{A}({M,N,\rho})}\left({A}-\Re\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2},\\ {\hat{A}}_{Q,k}^{(K)}=\arg\min_{{A}\in\mathcal{A}({M,N,\rho})}\left({A}-\Im\left\{{{Y}}_{k}^{(K)}\right\}/\alpha\right)^{2}.\end{cases}$ (25)
@@
    Apply the same CAF process at the transmitter to obtain the estimated in-band component, which is modeled as
    $\hat{{U}}_{k}^{(K)}=\alpha\hat{{X}}_{k}^{(K)}+\hat{{D}}_{k}^{(K)}.$ (26)
@@
    $=\alpha{X}_{k}+\left({D}_{k}-\hat{{D}}_{k}^{(K)}\right)+{W}_{k}.$ (28)
@@
Note that since our NENU constellations are designed based on a one-dimensional constellation, the complexity associated with the hard decision in (25) is considerably lower than that of a two-dimensional constellation, similar to the argument of demapper complexity discussed in Section III-E. In fact, if the modulation scheme is based on a general two-dimensional constellation denoted by $\mathcal{X}$, the hard decision must be performed in the complex plane as
${\hat{X}}_{k}^{(K)}=\arg\min_{{X}\in\mathcal{X}}\left|{X}-{{Y}}_{k}^{(K)}/\alpha\right|^{2},$ (29)
which requires comparing Euclidean distances among $\left|{\cal X}\right|=M^{2}$ symbols, instead of $\left|\mathcal{A}({M,N,\rho})\right|=N$ symbols. This observation also justifies the use of one-dimensional constellations for our BICM-OFDM systems employing CAF and CNC.
@@ IV-C Parameter Optimization for OFDM with CAF and CNC
The performance of OFDM with CAF and CNC significantly depends on system parameters, including the number of subcarriers, clipping ratio, the number of iterations, and the channel models [sun2021]. Therefore, it can only be evaluated by simulations for a given system setup. In what follows, we will first define the realistic system setup and then attempt to optimize shaping parameters based on computer simulations.
@@ IV-C1 System Setup
The specific system parameters as well as the channel model considered through the remainder of this section are listed in Table III: We consider a specific OFDM system with the number of subcarriers set to $N_{c}=1024$, with each subcarrier modulated by $32$-PAM per dimension, corresponding to $1024$-QAM per subcarrier. Channel coding is performed using an LDPC code compliant with the 5G New Radio (NR) standard, with a codeword length of $10240$ bits [etsi38138]. The code rate is fixed to $R_{c}=4/5$, which results in a spectral efficiency of $4.0$ bits per dimension, or $8.0$ bits per complex channel use. At the receiver, decoding is carried out using an offset min-sum decoding algorithm. Both AWGN and fading channels are considered. For the fading scenario, we adopt a frequency-selective Rayleigh fading channel following the 5G NR Tapped Delay Line (TDL)-B model [etsi2020138].
In the case of CAF and CNC operations, the oversampling factor is fixed to $J=4$ throughout this work. Based on the preliminary simulation with the above-mentioned system setup, we found that the clipping ratio of $\gamma_{\text{CR}}=1.5$ is the minimum acceptable value that does not cause the CNC operation to malfunction [^6] Optimal clipping ratio may be mathematically analyzed in terms of signal-to-noise plus distortion power ratio (SNDR) when CAF is applied [azolini], but since our system makes use of CNC at the receiver, its error rate performance depends on several parameters such as the number of subcarriers and constellations [sun2021]. As a result, it may become mathematically intractable. .
TABLE III: System parameters and channel model.
The number of subcarriers $N_{c}=1024$
Modulation $1024$-QAM ($32$-PAM each)
Clipping ratio $\gamma_{\text{CR}}=1.5$
Oversampling factor $J=4$
Maximum CNC iterations $K_{\text{max}}=10$
Channel code 5G NR LDPC [etsi38138]
Decoder Offset min-sum decoding
Code rate $4/5$
Codeword length 10240
Spectral efficiency 4.0 per dimension
Fading channel 5G NR Tapped Delay Line (TDL)-B [etsi2020138]
[…]
@@ IV-C2 Parameter Optimization
TABLE IV: System description and optimal parameters.
Mode CAF/CNC Target Channel Parameters $(N,\rho)$
A not applied AWGN/fading $(32,0.88)$
B applied AWGN $(24,0.96)$
C applied fading $(26,0.90)$
[…]
We optimize the NENU constellation for specific system configurations and channel conditions. We consider the following three system scenarios:
  • A.
    Without CAF, the system operates over either an AWGN channel or a fading channel.
  • B.
    With CAF and CNC, the system operates over an AWGN channel.
  • C.
    With CAF and CNC, the system operates over a fading channel.
In systems that utilize CAF with $\gamma_{\text{CR}}=1.5$, CNC is unable to completely eliminate clipping noise. Consequently, the effective channel becomes a mixture of AWGN and residual clipping noise, which may not possess a mathematically tractable representation. In subsequent studies focused on BMI evaluation and LLR computation, the noise variance $N_{0}$ specifically represents only the AWGN components. This limitation occurs because the residual clipping noise is affected by the constellation geometry and the transmitted symbol sequence, making it difficult to consistently incorporate it into $N_{0}$ across different modulation schemes.
Both the number of constellation points and their distribution should be designed to match the corresponding system condition. To this end, we evaluate the BER over a family of NENU constellations parameterized by $(N,\rho)$ and select the constellation that achieves a BER level of $10^{-4}$ at the lowest SNR as the representative design [^7] As discussed in Section III, the two parameters can also be selected based on the BMI. However, when the codeword length is not sufficient, and thus the gap between the channel capacity and the practical decoding waterfall region becomes non-negligible, BER-based optimization provides a more accurate criterion for parameter selection.. For each system scenario, the optimal parameter set $(N,\rho)$ obtained through preliminary simulations, as well as their descriptions, are listed in Table IV. In what follows, we refer to the resulting NENU constellations corresponding to the above systems A, B, and C as Modes A, B, and C, respectively, and denote them by $\mathcal{X}_{\sf A}$}, $\mathcal{X}_{\sf B}$}, and $\mathcal{X}_{\sf C}$}. Here, we note that in the absence of CAF, the optimal parameter set $(N,\rho)$ turned out to be identical for both AWGN and fading channels, and thus they are unified as Mode A. From the above results, we observe that the optimal number of constellation points depends on the channel conditions, which agrees with the observation in Section III. Without CAF, $N=M=32$ is selected as optimal due to the high spectral efficiency setup with $R_{c}=4/5$. In contrast, with CAF and CNC, smaller values of $N$ are preferred to better accommodate the residual nonlinear distortion that reduces the effective SNR, but the optimal values may differ depending on the channel models.
@@
It is worth mentioning that constellation design based on exhaustive numerical optimization algorithms, such as particle swarm optimization (PSO), does not rely on a predefined set of constellation candidates. Therefore, the optimized constellation, if found, should achieve near-optimal performance. However, if the objective function for optimization itself is computationally demanding, e.g., the BER after CAF and CNC processing, the exhaustive search should become challenging. On the other hand, due to limited degrees of freedom, the proposed shaping scheme requires evaluation over a tractable number of constellation candidates. It thus significantly reduces the computational burden and enables efficient optimization, even if the objective function is mathematically intractable and thus requires exhaustive simulations.
IV-C3 Total BMI for AWGN Channel
@@
Figure 9: Comparison of BMI values for 32-PAM constellations over an AWGN channel. The following three systems are evaluated: A1) Without CAF, A2) with CAF only, and A3) with CAF/CNC.
@@
Figure 10: Comparison of BMI values for 32-PAM constellations over a frequency-selective Rayleigh fading channel. The following two systems are evaluated: F1) Without CAF and F2) with CAF and CNC.
@@
For the purpose of investigating the effect of the parameter optimizations, we evaluate the total BMI over an AWGN channel, focusing on the following three systems:
@@
    BICM-OFDM without CAF (no PAPR reduction).
@@
    BICM-OFDM with CAF but without CNC.
@@
The results are compared in Fig. 9, where the NENU constellations are fixed as either $\mathcal{X}_{\sf A}$ or $\mathcal{X}_{\sf B}$ over the entire channel SNR. The uniform PAM is also shown as a reference. In A1, the constellation $\mathcal{X}_{\sf A}$ demonstrates superior performance compared to the others at the target information rate. On the other hand, in A2, the achievable rates of all constellations significantly deviate from the target as a result of severe clipping noise. Finally, in A3, the information rate is successfully recovered by CNC, allowing the target information rate to be achieved with a lower PAPR compared to A1. The constellation $\mathcal{X}_{\sf B}$ exhibits superior performance, yielding a greater gain over the constellation $\mathcal{X}_{\sf A}$. The results thus demonstrate the effectiveness of the proposed constellation design tailored to the BICM-OFDM system with CAF and CNC.
IV-C4 Total BMI for Fading Channel
We next consider the fading channel (5G NR TDL-B) for the following two systems:
@@ IV-C4 Total BMI for Fading Channel
    BICM-OFDM without CAF (no PAPR reduction).
@@
The BMI values are compared in Fig. 10, where the parameters of the NENU constellations are fixed as either $\mathcal{X}_{\sf A}$ or $\mathcal{X}_{\sf C}$ over the entire channel SNR, along with the uniform PAM as a reference. It is observed that for F1, the gains achieved by the proposed constellations slightly decrease compared to the results over the AWGN channel. This may be due to the fact that a single modulation scheme is applied across all subcarriers, while they experience different effective SNRs. Nevertheless, in F2, significant gains by the proposed shaping are still observed when compared at the target information rate. Additionally, the constellation $\mathcal{X}_{\sf C}$ (optimized for this system over a fading channel) offers enhanced performance over $\mathcal{X}_{\sf A}$.
@@ V Simulation Results
In the remainder of this paper, we present the system-level performance of the BICM-OFDM employing the proposed NENU constellations in the presence of CAF and CNC and make comparisons with other systems of similar transmitter complexity. The system parameters are chosen identically to Table III. As performance measures, we evaluate the distribution of PAPR as well as BER. We thus compare the following three systems:
@@
Note that S3 is another low-complexity PAPR reduction scheme comparable to CAF, and it is equivalent to a single-carrier system with a rectangular pulse shaping filter [myung2006single]. Therefore, its PAPR directly depends on the underlying constellations [ochiai_wcl2012]. For S3, a conventional minimum mean square error (MMSE) detector is employed at the receiver in the frequency-selective fading channel. For each system, the following modulation schemes are evaluated:
  • The standard uniform $1024$-QAM.
  • The proposed 1024-ary NENU constellations ($\mathcal{X}_{\sf A}$, $\mathcal{X}_{\sf B}$, or $\mathcal{X}_{\sf C}$) selected according to the system configuration.
  • The one-dimensional (1D) $1024$-NUC of ATSC 3.0.
  • The two-dimensional (2D) $1024$-NUC from [sillekens2022high].
The 1D-NUC and 2D-NUC are optimized according to their respective design algorithms for the same target spectral efficiency of $8.0$ bits per complex channel use and thus are suitable for the conventional OFDM system without CAF and CNC (i.e., S1 and S3).
@@ V-A CCDF of PAPR
Figure 11: Comparison of the CCDF of PAPR among the three systems: conventional OFDM without CAF, OFDM with CAF, and OFDM with DFT precoding (single-carrier), each employing the four different modulation schemes.
The CCDF curves of the PAPR among the three systems are shown in Fig. 11, where all the aforementioned $1024$-ary constellations are compared. We observe no difference among the four curves plotted for S1 and S2, suggesting that each subcarrier constellation does not affect PAPR distribution for OFDM. On the other hand, in S3, the selection of constellations significantly affects the dynamic range of the transmit signals, indicating that the use of constellation shaping increases the PAPR as expected [8227810], even though it is still less than the conventional OFDM (S1). In contrast, the results for S2 indicate that CAF effectively reduces the peak power of OFDM signals even in the presence of constellation shaping, achieving a PAPR reduction of $5.0$ dB compared to S1 when measured at a CCDF of $10^{-4}$.